OpenStudy (anonymous):

How to add 1/(-6)-1/4?

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \bf \frac{1}{-6}- \frac{1}{4} }\) this ?

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \bf \frac{a}{-b}= -\frac{a}{b} }\) do this and then find least common denominator and add.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can u shoe me?

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \bf \frac{1}{-6} -\frac{1}{4} }\) \(\Huge\color{blue}{ \bf -\frac{1}{6} -\frac{1}{4} }\) \(\Huge\color{blue}{ \bf -\frac{4}{24} -\frac{6}{24} }\) \(\Huge\color{blue}{ \bf -\frac{10}{24} }\)